# Eliciting and Distinguishing Between Weak and Incomplete Preferences: Theory, Experiment and Computation

### v5

@@ 2.2         Distinguishing Between Indifference and Incomparability/Indecisiveness
indecisive between $x$ and $y$ if and only if $x,y\in A$Rightarrow $x,y\not\in C(A)$.
@@ 2.3         Measuring a Choice Correspondence’s Proximity to A Model:
Recall that a dataset $\mathcal{D}$ consists of a finite collection of observations $\big{(}A,C(A)\bigr{)}$, where $\emptyset\neq A\subseteq X$ and $\emptyset\subseteq C(A)\subseteq A$. Denote by $\mathbf{D}$ the collection of all such datasets. In the important special case where $|C(A)|=1$ is true for every pair $\big{(}A,C(A)\bigr{)}$ in $\mathcal{D}\in\mathbf{D}$ one can measure the proximity of $\mathcal{D}$ to the model of utility maximization with strict preferences with the classic HM index of rationality due to Houtman and Maks (1985) [^3] An axiomatization of this index in a rich dataset environment is given in Apesteguia and Ballester (2015)., defined by the function $\mathtt{HM}:\mathbf{D}\rightarrow\mathbb{N}$ such that
$\mathtt{HM}(\mathcal{D})$ ::=:= $\min_{\succ\in\mathcal{P}}\left|\big{(}A,C(A)\bigr{)}\in\mathcal{D}:C(A)\neq C_{\succ}(A)\right|,$ (3)
@@
For this reason, instead of generalizing (3) to the multi-valued choice case in what appears to be a counter-intuitively punitive way, we propose a generalization that explicitly accounts for the content-similarity of sets $C(A)$ and $C_{\succsim}(A)$ across all pairs $\big{(}A,C(A)\bigr{)}\in\mathcal{D}$ and every weak preference relation $\succsim\in\mathcal{R}$. More specifically, to assess how dissimilar is the observed choice $C(A)$ to the choice $C_{\succsim}(A)$ that is optimal under some $\succsim\in\mathcal{R}$ we use the classic Jaccard dissimilarity metric (Jaccard, 1901), whereby
$J\big{(}C(A),C_{\succsim}(A)\bigr{)}$ ::=:= $\dfrac{|C(A)\cup C_{\succsim}(A)|-|C(A)\cap C_{\succsim}(A)|}{|C(A)\cup C_{\succsim}(A)|}$ (4)
@@
This function, which is widely applied in the life sciences and elsewhere, is a proper metric (Marczewski and Steinhaus, 1958; Levandowsky and Winter, 1971; Gerasimou, 2024) and identifies the dissimilarity between two sets as the proportion of their non-common elements relative to the total number of distinct elements between them. With this definition in place, we can now introduce the Jaccard-Houtman-Maks (JHM) generalization of (3) for the case of possibly multi-valued choice data as the function $\mathtt{JHM}:\mathbf{D}\rightarrow\mathbb{Q}_{+}$ where
$\mathtt{JHM}(\mathcal{D})$ ::=:= $\min_{\succsim\in\mathcal{R}}\sum_{\big{(}A,C(A)\bigr{)}\in\mathcal{D}}J\big{(}C(A),C_{\succsim}(A)\bigr{)}$ (6)
@@ 3.1         General Remarks
There are several reasons why the experimental choice alternatives were selected to be gift-card pairs, and those ones in particular. First, gift cards can be thought of as restricted forms of cash that can be used for consumption, informally traded or, indeed, gifted. As such, they are intrinsically valuable. Second, these particular gift cards were issued by some of the most popular leisure and grocery destinations for the local student population, and were all within a short walking distance from each other. Hence, all were expected to be desirable to everyone in this experiment’s subject pool. Finally, presenting the choice alternatives as gift card pairs rather than as separate gift cards is realistic (many online retailers invite consumers to choose their “gift card bundle”) and could potentially lead to some relatively hard decisions.
Out of the 63 non-empty menus that are derivable from this set of 6 alternatives, subjects were presented with the 15 binary, 20 ternary and 15 quaternary menus. Each of the 50 distinct menus was presented once and the order of menu presentation was randomized and differed between subjects. This choice domain is therefore rich, heterogeneous and symmetric in the sense that all alternatives are feasible in exactly the same number of menus (see also Appendix B for more details on this). Furthermore, the total number of 50 decisions coincides with the one used in experiments on budget-constrained choice from Arrow-Debreu securities such as Choi, Fisman, Gale, and Kariv (2007) and Halevy, Persitz, and Zrill (2018). Unlike these studies, the choice objects in this experiment were riskless and presented naturalistically, while subjects did not operate under a budget constraint.
The experiment’s computer interface was programmed in Qualtrics and executed in full-screen mode on a web browser that prevented subjects from exiting the interface without the experimenter’s intervention. Subject recruitment was done with ORSEE (Greiner, 2015). All menus appeared as unnumbered vertical choice lists. The “I’m not choosing now” option in the Non-Forced-Choice treatment was always the last item. Because subjects often had to scroll down to find and select that option, this positioning means that it was often physically harder for them to avoid/delay choice. There were a total of 141 participants in each of the two treatments. The 9 subjects who always deferred or always chose everything were excluded from the analysis because their choice behaviour is completely uninformative. Every subject received both their gift-card and cash rewards, as explained below. A nine-question understanding quiz preceded the main part of the experiment in each of the two treatments. Subjects could not proceed to the main part of the experiment until they answered all questions correctly.
@@ 3.4         Payment
As soon as subjects finished all tasks, their randomly selected menu showed up on their screens, together with the reminder of the decision they had made at this menu. As an additional incentive for subjects to make deliberated and non-rushed decisions, they were told from the beginning that no participant would be able to receive their rewards and leave the lab in the first 50 minutes of the session. The experimenter (this author) went to each subject’s desk once they were finished (and after this threshold was exceeded), asked them about their final choice at this menu, and later gave them their cash and gift card rewards accordingly. Subjects who had chosen everything at their randomly selected menu were in turn invited to the experimenter’s desk and, after the appropriate numerical range was specified on the random-number generating website https://random.org, and the way in which the numbers in this range were mapped to the relevant gift card pairs was agreed, a random number was generated to determine the pair they would be rewarded with. The total cost from payments to experimental subjects was approximately £6,270.
@@ 3.5         Discussion
Table 3: Incentives in the Forced- and Free-Choice treatments are structured around the subjects’ 1st and 2nd decisions at their randomly selected menu (denoted here by $A$).
1st decision 2nd decision Choice Cash Possible (main stage) (payoff stage) reward reward interpretation (a) Forced-Choice Treatment $C^{1}(A)=A$ none random initial Revealed total indifference possible $a\in A$ endowment $I$ is costless & responded to literally $C^{1}(A)\subset A$ $C^{2}(A)\in C^{1}(A)$ $C^{2}(A)$ $I$ Stable revealed preference is costless $C^{1}(A)\subset A$ $C^{2}(A)\not\in C^{1}(A)$ $C^{2}(A)$ $I_{r}<I$ Unstable revealed preference is costly (b) Free-Choice Treatment $C^{1}(A)=A$ none random initial Revealed total indifference possible $a\in A$ endowment $I$ is costless & responded to literally $\emptyset\neq C^{1}(A)\neq A$ $C^{2}(A)\in C^{1}(A)$ $C^{2}(A)$ $I$ Stable revealed preference is costless $\emptyset\neq C^{1}(A)\neq A$ $C^{2}(A)\not\in C^{1}(A)$ $C^{2}(A)$ $I_{r}<I$ Unstable revealed preference is costly, and more so than $C^{1}(A)=\emptyset$ $C^{2}(A)\in A$ $C^{2}(A)$ $I_{d}$ :  $I_{r}<I_{d}<I$ revealed indecisiveness
@@
At the same time, if a utility-maximizing subject has one or more optimal alternative(s) at the menu, and these are strictly better than something else feasible, then choosing everything is dis-incentivized by the design because doing so comes with the risk of potentially receiving an inferior alternative as a reward. The design instead incentivizes such a subject to select those and only those options they prefer the most, as this enables them to choose one of these superior options at the end and to also receive their full cash endowment. For a utility-maximizing agent, therefore, the design in this treatment novelly combines incentive-compatibility for truthful preference revelation in the sense of Azrieli, Chambers, and Healy (2018) with a standard interpretation of multi-valued choice that is articulated in Kreps (2012, p.2) as follows: “The story is that the consumer chooses one element of $A$. Nonetheless, we think of $c(A)$ as a subset of $A$, not a member or element of $A$. This allows for the possibility that the consumer is happy with any one of the several elements of $A$, in which case $c(A)$ lists all those elements. When she makes a definite choice of a single element, say $x$, out of $A$–when she says in effect, ‘I want $x$ and nothing else’–we write $c(A)=\{x\}$, or the singleton set consisting of the single element $\{x\}$. But if she says, ‘I would be happy with either $x$ or $y$’, then $c(A)=\{x,y\}$.”
@@
The choice correspondence $C^{1}(\cdot)$ for a Free-Choice subject on the other hand is possibly empty- and multi-valued; hence, it satisfies $\emptyset\subseteq C^{1}(A)\subseteq A$ for every menu $A$. As in the Forced-Choice treatment, however, $C^{2}(\cdot)$ assigns a single chosen alternative to the relevant subject’s randomly selected menu. In this treatment too, moreover, a utility-maximizing Free-Choice subject is (weakly) incentivized to choose their most preferred option(s) at every menu, with the opportunity to defer at a cost being irrelevant to them. Now suppose again that the subject is not a utility maximizer but has incomplete preferences and cannot compare any two of the feasible options at a menu. Then, depending on the individual’s subjective perception of the decision’s importance, they may either opt to choose everything and end up with something at random or opt instead to incur the relatively small cost $I-I_{def}$ to delay making an active choice at that menu themselves. The former kind of behaviour could be seen as analogous to recent findings suggesting that decision makers prefer to randomize when they are faced with a difficult decision repeatedly (Agranov and Ortoleva, 2017; Dwenger, Kübler, and Weizsäcker, 2018). Importantly though, it could also be compatible with the model of undominated choice with incomplete preferences –discussed in Section 4– whenever every feasible option is incomparable to all others. By contrast, deferring at a small cost could be seen as a manifestation of indecisiveness-driven deferral (Tversky and Shafir, 1992; Danan and Ziegelmeyer, 2006; Costa-Gomes, Cueva, Gerasimou, and Tejiščák, 2022) and may potentially be compatible with the model of dominant choice with incomplete preferences that was discussed in the previous section.
3.6         A Potential Criticism
@@
It is worth commenting, finally, on the discrepancy in the way that the design deals with situations where $C(A)=A$ or $\emptyset\neq C(A)\subset A$ at the payoff-relevant menu $A$. Recall that in the former case subjects receive their reward from $C(A)$ at random and a full cash reward alongside it, whereas in the latter case they can themselves choose anything from $A$, possibly at the cost of receiving the reduced cash reward if their two choices at $A$ exhibit a reversal. An alternative design choice here would be to break this dichotomy by applying the uniform randomization rule at $C(A)$ in both situations. This would have been a perfectly legitimate way to proceed in our environment of multi-valued choice. Introducing this dichotomy here was driven by the motivation to have a design that remained comparable to the one in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). The one proposed here achieves this goal because it does indeed reduce to the one proposed by these authors in the special case where $C(A)$ is required to be single-valued. With the alternative design mentioned above there would be no role for the reduced cash payment $I_{rev}$ that is associated with a choice reversal. The presence of this additional payment parameter, however, and its lower value compared to that of the full-payment parameter $I$, is potentially useful in instilling a mindset towards preference-guided choices by subjects. On the other hand, the unified choice-reward rule where randomization is applied on any non-empty $C(A)$ is immune to the possibility of subjects’ choosing dominated alternatives, which in theory is left open by the dichotomous rule in the current design. While this theoretical possibility can indeed not be ruled out, it is worth keeping in mind that a subject who includes dominated alternatives in their first-round choice $C(A)\subset A$ not only does not benefit from doing so but, in fact, spends additional time and effort –thereby incurring a cost– in the process.

### v6

@@ 2.2   Distinguishing Between Indifference and Incomparability/Indecisiveness
indecisive between $x$ and $y$ if and only if $x,y\in A$ $\Rightarrow$ $x,y\not\in C(A)$.
@@ 2.3   Measuring a Choice Correspondence’s Proximity to A Model:
Our non-parametric model-fitting analysis extends the classic Houtman-Maks (HM) (Houtman and Maks, 1985) method that was outlined in the introduction by accounting for: (i) multi-valued (as well as empty-valued) choice data, and (ii) the possibility that the decision maker generating such data behaves under a model that is more general than utility maximization. Indeed, by finding which model is closest to explaining each subject’s behaviour, this method effectively allows for recovering –possibly approximately, and with the standard as if qualifications in place– both the individual’s deterministic decision rule and their preferences conditional on that rule. This explains the “model-based” part of the claimed extension. In order to also account for the general multi-valuedness of subjects’ choices and of the model-predicted optimal choices, our proposed extension of the HM method computes the (dis-)similarity between these sets –which intuitively captures the proximity of observed and model-predicted choices– by means of the classic Jaccard (dis-)similarity metric between finite sets of discrete objects (Jaccard, 1901; Marczewski and Steinhaus, 1958; Levandowsky and Winter, 1971). This, in particular, allows one to have a more nuanced understanding of when a deviation of a subject’s set of chosen alternatives from the model-predicted ones is a “severe” mistake or a relatively “mild” one, contrary to the binary “pass/fail” approach of the straightforward extension of the HM method in this environment. Hence, we will refer to this as the Jaccard-Houtman-Maks (JHM) method.
More formally, recall that a dataset $\mathcal{D}$ consists of a finite collection of observations $\big{(}A,C(A)\bigr{)}$, where $\emptyset\neq A\subseteq X$ and $\emptyset\subseteq C(A)\subseteq A$. Denote by $\mathbf{D}$ the collection of all such datasets. In the important special case where $|C(A)|=1$ is true for every pair $\big{(}A,C(A)\bigr{)}$ in $\mathcal{D}\in\mathbf{D}$ one can measure the proximity of $\mathcal{D}$ to the model of utility maximization with strict preferences with the classic HM index of rationality due to Houtman and Maks (1985) [^5] An axiomatization of this index in a rich-dataset environment is given in Apesteguia and Ballester (2015)., defined by the function $\mathtt{HM}:\mathbf{D}\rightarrow\mathbb{N}_{0}$ such that
$\mathtt{HM}(\mathcal{D})$ $:=$ $\min_{\succ\in\mathcal{P}}\left|\big{(}A,C(A)\bigr{)}\in\mathcal{D}:C(A)\neq C_{\succ}(A)\right|,$ (3)
@@
[…]
$J\big{(}C(A),C_{\succsim}(A)\bigr{)}$ $:=$ $\dfrac{|C(A)\cup C_{\succsim}(A)|-|C(A)\cap C_{\succsim}(A)|}{|C(A)\cup C_{\succsim}(A)|}$ (4)
@@
This function, which is widely applied in the life sciences and elsewhere, is a proper metric (Marczewski and Steinhaus, 1958; Levandowsky and Winter, 1971; Gerasimou, 2024) and identifies the dissimilarity between two sets as the proportion of their non-common elements relative to the total number of distinct elements between them. With this definition in place, we can now formally introduce our proposed generalization of (3) for the case of possibly multi-valued choice data as the function $\mathtt{JHM}:\mathbf{D}\rightarrow\mathbb{Q}_{+}$ where
$\mathtt{JHM}(\mathcal{D})$ $:=$ $\min_{\succsim\in\mathcal{R}}\sum_{\big{(}A,C(A)\bigr{)}\in\mathcal{D}}J\big{(}C(A),C_{\succsim}(A)\bigr{)}$ (6)
@@ 3.1   General Remarks
There are several reasons why the experimental choice alternatives were selected to be gift-card pairs, and those ones in particular. First, gift cards can be thought of as restricted forms of cash that can be used for consumption, informally traded or, indeed, gifted. As such, they are intrinsically valuable. Second, these particular gift cards were issued by some of the most popular leisure and grocery destinations for the local student population, and were all within a short walking distance from each other. Hence, all were expected to be desirable to everyone in this experiment’s subject pool. Finally, presenting the choice alternatives as gift card pairs proxies a realistic situation (many online retailers invite consumers to choose between multi-store gift card bundles rather than individual-store gift cards) and could potentially lead to some relatively hard decisions.
Out of the 63 non-empty menus that are derivable from this set of 6 alternatives, subjects were presented with the 15 binary, 20 ternary and 15 quaternary menus. Each of the 50 distinct menus was presented once and the order of menu presentation was randomized and differed between subjects. This choice domain is therefore rich, heterogeneous and symmetric in the sense that all alternatives are feasible in exactly the same number of menus (see also Appendix B for more details on this).
The experiment’s computer interface was programmed in Qualtrics and executed in full-screen mode on a web browser that prevented subjects from exiting the interface without the experimenter’s intervention. Subject recruitment was done with ORSEE (Greiner, 2015). All menus appeared as unnumbered vertical choice lists. The “I’m not choosing now” option in the Free-Choice treatment was always the last item. Because subjects often had to scroll down to find and select that option, this positioning means that it was often physically harder for them to avoid/delay choice. There were a total of 141 participants in each of the two treatments. The 9 subjects who always deferred or always chose everything were excluded from the analysis because their choice behaviour is completely uninformative. Every subject received both their gift-card and cash rewards, as explained below. A nine-question understanding quiz preceded the main part of the experiment in each of the two treatments. Subjects could not proceed to the main part of the experiment until they answered all questions correctly.
@@ 3.4   Payment
As soon as subjects finished all tasks, their randomly selected menu showed up on their screens, together with the reminder of the decision they had made at this menu. As an additional incentive for subjects to make deliberated and non-rushed decisions, they were told from the beginning that no participant would be able to receive their rewards and leave the lab in the first 50 minutes of the session. The experimenter (this author) went to each subject’s desk once they were finished (and after this threshold was exceeded), asked them about their final choice at this menu, and later gave them their cash and gift card rewards accordingly. Subjects who had chosen everything at their randomly selected menu were in turn invited to the experimenter’s desk and, after the appropriate numerical range was specified on the random-number generating website https://random.org, and the way in which the numbers in this range were mapped to the relevant gift card pairs was agreed, a random number was generated to determine the pair they would be rewarded with. The total payment per subject was approximately £22.2.
@@ 3.5   Discussion
Table 3: Incentives in the Forced- and Free-Choice treatments are structured around the subjects’ 1st and 2nd decisions at their randomly selected menu (denoted here by AA).
1st decision 2nd decision Choice Cash Possible (main stage) (payoff stage) reward reward interpretation (a) Forced-Choice Treatment $C^{1}(A)=A$ none random initial Revealed total indifference possible $a\in A$ endowment $I$ is costless & responded to literally $C^{1}(A)\subset A$ $C^{2}(A)\in C^{1}(A)$ $C^{2}(A)$ II Stable revealed preference is costless $C^{1}(A)\subset A$ $C^{2}(A)\not\in C^{1}(A)$ $C^{2}(A)$ $I_{r}<I$ Unstable revealed preference is costly (b) Free-Choice Treatment $C^{1}(A)=A$ none random initial Revealed total indifference possible $a\in A$ endowment $I$ is costless & responded to literally $\emptyset\neq C^{1}(A)\neq A$ $C^{2}(A)\in C^{1}(A)$ $C^{2}(A)$ II Stable revealed preference is costless $\emptyset\neq C^{1}(A)\neq A$ $C^{2}(A)\not\in C^{1}(A)$ $C^{2}(A)$ $I_{r}<I$ Unstable revealed preference is costly, and more so than $C^{1}(A)=\emptyset$ $C^{2}(A)\in A$ $C^{2}(A)$ $I_{d}$ :  $I_{r}<I_{d}<I$ revealed indecisiveness
@@
At the same time, if a utility-maximizing subject has one or more optimal alternative(s) at the menu, and these are strictly better than something else feasible, then choosing everything is dis-incentivized by the design because doing so comes with the risk of potentially receiving an inferior alternative as a reward. The design instead incentivizes such a subject to select those and only those options they prefer the most, as this enables them to choose one of these superior options at the end and to also receive their full cash endowment. For a utility-maximizing agent, therefore, the design in this treatment combines incentive-compatibility for truthful preference revelation in the sense of Azrieli, Chambers, and Healy (2018) with a standard interpretation of multi-valued choice that is articulated in Kreps (2012, p.2) as follows: “The story is that the consumer chooses one element of $A$. Nonetheless, we think of $c(A)$ as a subset of $A$, not a member or element of $A$. This allows for the possibility that the consumer is happy with any one of the several elements of $A$, in which case $c(A)$ lists all those elements. When she makes a definite choice of a single element, say $x$, out of AA–when she says in effect, ‘I want $x$ and nothing else’–we write $c(A)=\{x\}$, or the singleton set consisting of the single element $\{x\}$. But if she says, ‘I would be happy with either $x$ or $y$’, then $c(A)=\{x,y\}$.”
@@
The choice correspondence $C^{1}(\cdot)$ for a Free-Choice subject on the other hand is possibly empty- and multi-valued; hence, it satisfies $\emptyset\subseteq C^{1}(A)\subseteq A$ for every menu $A$. As in the Forced-Choice treatment, however, $C^{2}(\cdot)$ assigns a single chosen alternative to the relevant subject’s randomly selected menu. In this treatment too, moreover, a utility-maximizing Free-Choice subject is (weakly) incentivized to choose their most preferred option(s) at every menu, with the opportunity to defer at a cost being irrelevant to them. Now suppose again that the subject is not a utility maximizer but has incomplete preferences and cannot compare any two of the feasible options at a menu. Then, depending on the individual’s subjective perception of the decision’s importance, they may either opt to choose everything and end up with something at random or opt instead to incur the relatively small cost $I-I_{def}$ to delay making an active choice at that menu themselves. The former kind of behaviour could be seen as analogous to recent findings suggesting that decision makers prefer to randomize when they are faced with a difficult decision repeatedly (Agranov and Ortoleva, 2017; Dwenger, Kübler, and Weizsäcker, 2018); we discuss this possibility in Appendix C.2. Importantly though, it could also be compatible with the model of Undominated Choice with Incomplete Preferences whenever every feasible option is incomparable to all others. By contrast, deferring at a small cost could be seen as a manifestation of indecisiveness-driven deferral (Tversky and Shafir, 1992; Danan and Ziegelmeyer, 2006; Costa-Gomes, Cueva, Gerasimou, and Tejiščák, 2022) and may potentially be compatible with the model of Dominant Choice with Incomplete Preferences.
3.6   Remarks on the Asymmetric Incentives in the Design
@@
It is worth commenting on the discrepancy in the way that the design deals with situations where $C(A)=A$ or $\emptyset\neq C(A)\subset A$ at the payoff-relevant menu $A$. Recall that in the former case subjects receive their reward from $C(A)$ at random and a full cash reward alongside it, whereas in the latter case they can themselves choose anything from $A$, possibly at the cost of receiving the reduced cash reward if their two choices at $A$ exhibit a reversal. An alternative design choice here would be to break this dichotomy by applying the uniform randomization rule at $C(A)$ in both situations. This would have been a perfectly legitimate way to proceed in our environment of multi-valued choice. Introducing this dichotomy here was driven by the motivation to have a design that remained comparable to the one in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). The one proposed here achieves this goal because it does indeed reduce to the one proposed by these authors in the special case where $C(A)$ is required to be single-valued. With the alternative design mentioned above there would be no role for the reduced cash payment $I_{rev}$ that is associated with a choice reversal. The presence of this additional payment parameter, however, and its lower value compared to that of the full-payment parameter $I$, is potentially useful in instilling a mindset towards preference-guided choices by subjects. On the other hand, the unified choice-reward rule where randomization is applied on any non-empty $C(A)$ is immune to the possibility of subjects’ choosing dominated alternatives, which in theory is left open by the dichotomous rule in the current design. While this theoretical possibility can indeed not be ruled out, it is worth keeping in mind that a subject who includes dominated alternatives in their first-round choice $C(A)\subset A$ not only does not benefit from doing so but, in fact, spends additional time and effort –thereby incurring a cost– in the process.
4   Individual-Level Analysis
4.1   Non-Parametric Model Fitting
We now turn to the tests of our possibly multi- and/or empty-valued experimental choice data for compatibility with each of the three models introduced in Section 2. These tests were done by applying the JHM method that was introduced in Section 2.3. We refer to the JHM score of a model on a given subject’s dataset as the model’s distance score on that dataset. [^6] We use the term “distance score” rather than “distance” partly because the underlying function is not a proper metric and partly to highlight the ranking aspect of this JHM-based goodness-of-fit approach. The model-based extension of HM was for single-valued (forced or free) choices was originally introduced in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016, 2022) [^7] The 2016 working-paper version of that study indirectly constructed choice correspondences by augmenting the subjects’ actual choices with their survey responses in indifference-vs-preference questions that followed binary menus only. Before constructing the subjects’ indifference classes from these data, the survey responses were analysed for consistency, and responses that eventually led to intransitive indifferences were discarded. Whenever this was the case, the subjects’ actual choices were interpreted as revealing strict preferences, regardless of the survey response. Since Gerasimou and Tejiščák (2018), moreover, this has been extended considerably and made freely available online in the open-source desktop application [Uncaptioned image]. [^8] With retainment of its original focus on the model of Rational Choice/Utility Maximization with strict preferences, the HM approach has been extended in other intuitive directions by Apesteguia and Ballester (2015) and Dean and Martin (2016). However, although the “Swaps” measure of Apesteguia and Ballester (2015) that pertains to general choice environments is also readily computable in [Uncaptioned image], it is currently unclear how it should be computed when subjects choose multiple items from the same menu, or how it should be extended in models that generalise utility maximization.
More specifically, a brute-force algorithm was implemented for these combinatorial-optimi-zation computations. This involved the production of all choice datasets that are generated by all possible instances of every model, and comparing each such dataset against every subject’s own dataset in order to find the model(s) and instance(s) that are closest to it in the minimum distance-score sense. This method therefore detects perfect as well as approximate model fits, and in the latter case it quantifies the approximation in an intuitive way that may be interpreted as the number of “mistakes” made by an agent who might be portrayed as following a particular decision rule. [^9] For other approximations that have been proposed or applied recently for/at distinct classes of models and choice environments we refer the reader to Choi, Fisman, Gale, and Kariv (2007); Echenique, Lee, and Shum (2011); Choi, Kariv, Müller, and Silverman (2014); Apesteguia and Ballester (2015); Dean and Martin (2016); Bouacida and Martin (2021); Dembo, Kariv, Polisson, and Quah (2021); Fudenberg, Liang, and Gao (2023); Cherchye, Demuynck, Lanier, and Rock (2023); Echenique, Imai, and Saito (2023) and references therein.
Despite the large numbers of weak orders (4,683), partial orders (130,023) and, especially, incomplete preorders (209,527) that are defined on a set of 6 alternatives (OEIS, 2021), the above tool makes such an exact computation possible very quickly –in less than 12 seconds– for all subjects in each treatment, for each of the three models. Furthermore, although the brute-force algorithm is linear in the number of subjects but exponential in the number of alternatives, the model-rich distance-score method itself is scalable and can be extended to analyse datasets that are derived from much larger sets of alternatives by employing powerful open-source constraint solvers.
Undominated Dominant Utility Choice with Choice with Maximization Incomplete Incomplete All Preferences Preferences Forced-Choice treatment ($N=138$) Subjects with JHM = 0 5 (4%) 0 5 (4%) Subjects with JHM $\leq$ 5 61 (44%) 16 (11.5%) 77 (56%) Mean/median best score (JHM $\leq$ 5) 2.01/1.83 3.56/4 2.34/2 Mean / median best-model 1.23/1 1.19/1 1.22/1 preference orderings (JHM $\leq$ 5) Minimum score in simulations 14.87 13.92 Free-Choice treatment ($N=135$) Subjects with JHM = 0 3 (2%) 0 7 (5%) 10 (7%) Subjects with JHM $\leq$ 5 30 (22%) 7 (5%) 37 (27%) 74 (55%) Mean/median best score (JHM $\leq$ 5) 2.12/1.75 2.37/2.08 1.93/2 2.05/2 Mean/median best-model 1/1 1/1 1.16/1 1.08/1 preference orderings (JHM $\leq$ 5) Minimum score in simulations 21.03 19.67 20.08
(a) Forced-Choice treatment
(b) Free-Choice treatment
For human subjects, Rational Choice often tied with one, but never both, of the other two models. The classification presented in Table 4 broke ties in favour of that model whenever a subject’s JHM distance score was produced by Rational Choice and another model (see the notes of Table 4 for more details). Under this tie-breaking assumption, and in line also with model-approximation methods used in the above-cited studies, the two exhibits show the number, proportion and relative-frequency distributions of distinct subjects that are on average within 10% (equivalently, within 5 decisions) away from being explainable perfectly by an instance of some model, separately for each of the two experimental treatments. We chose this conservative approximation range for two reasons: (i) simulations suggest that a JHM distance score of 5 for any of the three models is extremely unlikely to occur randomly in this decision environment (see Table 4 and Figure 1); (ii) for subjects with a JHM score not greater than 5 there is typically only one best-matching preference ordering that explains their behaviour under the respective subject-optimal model.
The proportion of Forced-Choice subjects with a perfect or 10%-approximate model fit was 4% (all under Rational Choice) and 56%, respectively, with 44% and 11.5% of those in the latter group classified under Rational Choice and Undominated Choice with Incomplete Preferences. Furthermore, the proportion of Free-Choice subjects with a perfect or 10%-approximate fit was 7% and 55%, respectively. 22%, 5% and 27% of all subjects in this treatment were categorized under Rational Choice, Undominated and Dominant Choice with Incomplete Preferences, respectively. In addition, 3 and 7 of these subjects were perfectly compatible with the first and third model, respectively, with one subject among the former three and five among the latter seven revealing indifferences in their weakly ordered and weakly preordered preferences.
The model fit is slightly better on average in the Free-Choice treatment than in the Forced-Choice one (mean/median best JHM scores: 2.05/2 vs 2.34/2), although the distributions are not significantly different. Similarly, and in line with the results from the consistency analysis that is presented in Section 5, the proportion of subjects who were best-matched by one of the two models of consistent active choices (i.e. Rational Choice and Dominant Choice with Incomplete Preferences) is higher in the Free-Choice treatment (49% vs 44%), while the proportion of subjects best-matched by the third model (Undominated Choice with Incomplete Preferences) that predicts generally inconsistent active choices is lower in that treatment (5% vs 11.5%). Moreover, when Rational Choice was not the best-matching model, its distance score was on average 1.4 and 5.7 units higher in the Forced- and Free-Choice treatments, respectively. Finally, the highest JHM distance score of 5 for experimental subjects in this approximate model-fitting analysis is far below the corresponding minimum scores of simulated random-behaving subjects, which ranged between 14 and 21.
Figure 2: All 37 subjects that were best-matched by Dominant Choice with Incomplete Preferences were unlikely to have achieved this classification at random conditional on the menus where they deferred.
For the 37 subjects who are best-explained by Dominant Choice with Incomplete Preferences, moreover, we conducted an additional robustness check. By analogy to the way in which the “predictive success” (Selten, 1991) score of the Rational-Choice model on Free-Choice subjects’ active-choice consistency is evaluated in Section 5.4 below, we generated a set of 10,000 random-behaving simulated subjects for each of the 37 human experimental participants who were best-matched by that model, holding fixed in each such subject’s simulations block the menus at which the relevant human participant deferred. This allows for evaluating the predictive success of the Dominant-Choice model’s 10%-approximation range in this decision environment. Figure 2 shows the histogram with JHM distance-score differences for this model between the 2.5th percentile value in each subject-specific 10k-simulations block and the relevant subject’s actual score. For all 37 subjects the difference is positive and typically much greater than 10 score units (i.e. active-choice or deferral decisions). This analysis suggests that Dominant Choice with Incomplete Preferences is not excessively permissive in this environment, and that its good fit is useful towards explaining behaviour and eliciting preferences.
As was highlighted earlier, the collected choice data and non-parametric combinatorial-optimization method presented earlier jointly allow us to separate the strict part from the potential indifference and/or incomparability parts of each subject’s revealed preferences, conditional on the model that best explains the subject’s overall behaviour. For the vast majority of all subjects that are perfectly or approximately matched by some model this analysis reveals a unique weak/strict and complete/incomplete preference ordering that generates their minimum JHM distance score, with the mean (median) number of such orderings being 1.22 (1) and 1.08 (1) in the Forced- and Free-Choice treatments. Thus, despite the relatively large number of 50 distinct menus and decisions in the experiment and the fact that this large number nevertheless leads to an incomplete collection of menus, this model-rich analysis generates sharp subject-specific preference estimates.
Figure 3: A subject’s complete weak preferences, optimally recovered by Utility Maximization.
Figure 4: A subject’s incomplete preferences, optimally recovered by Undominated Choice with Incomplete Preferences, with (left) or without (right) indifferences.
Refer to captionRefer to caption
Figure 5: A subject’s incomplete weak preferences, optimally recovered by Dominant Choice with Incomplete Preferences.
We now turn to the important task of distinguishing between indifference and incomparability. To this end, it is perhaps worth recalling first that indifferences are generically non-existent in a large class of the benchmark Bewley (2002) class of incomplete preference relations under uncertainty that are defined on a space of bundles of continuous commodities or uncertain acts (Gerasimou, 2018b). By contrast, however, this is very much not the case when such preferences are over finitely many discrete options, as in the choice environment analysed in this paper. Indeed, our analysis uncovers that out of the 60 subjects (22% of the total) that the two models of incomplete-preference maximization explain optimally across the two treatments, for 46 of them (77%) it is possible to separate the strict, indifference and incomparability/indecisiveness components of their preferences (robustly so for the model of Dominant Choice with Incomplete Preferences). The empirical documentation of this theory-based separation of the three binary relations that are possible when preferences are incomplete is a novel contribution of this paper. Figures 4 and 5 show the directed graphs of two subjects’ incomplete weak/strict preferences that were optimally recovered by these models; Figure 3 does so for a utility-maximizing subject.
Table 5: Revealed indifferences conditional on each subject’s best-fitting model.
Indifference-revealing Mean (median)
subjects indifferences
Utility Maximization   71 (79%)  2.89 (2)
Dominant Choice with Incomplete Preferences   28 (76%)  2.26 (2)
Undominated Choice with Incomplete Preferences   16 (69.5%)  0.69 (0.5)
Overall   115 (77%)  2.4 (1.08)
Looking across the three models and two treatments, Table 5 summarizes the key revealed-indifference information that is derived from each subject’s best-fitting model and demonstrates the descriptive relevance of indifference in these data. In particular, 79%, 76% and 69.5% of all subjects who are best-explained by Utility Maximization, Dominant Choice and Undominated Choice with Incomplete Preferences, respectively, have at least one model-compatible preference relation that reveals at least one indifference between distinct gift-card pairs. Across all models, subjects and their respective model-compatible preference relations, finally, the average (median) number of indifference comparisons between distinct gift-card pairs is 2.4 (1.08). Considering that there are 15 possible binary comparisons on a set of 6 elements, this means that the indifference relation accounts for 16% of subjects’ preference comparisons, on average. This proportion is highest for Utility Maximization (19%), followed by Dominant (15%) and Undominated Choice with Incomplete Preferences (5%).
4.3   Cross-Validation Test of Model-Fitting Analysis